OpenStudy (anonymous):

Algebra 1 help - Geometry/Algebraic expressions

jimthompson5910 (jim_thompson5910):

I'm getting the same for both as well

OpenStudy (anonymous):

O so I got those questions right?

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thank you! I have a few more that use this same graph...can you check my answers on those please? Hold on, I will post it.....

jimthompson5910 (jim_thompson5910):

3 is incorrect

jimthompson5910 (jim_thompson5910):

how many students chose skateboarding?

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

how many total?

jimthompson5910 (jim_thompson5910):

3/20 = ???

OpenStudy (anonymous):

i am sort of confused....what do you do then? Divide?

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

convert that fraction to decimal form

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

0.15 = 15%

jimthompson5910 (jim_thompson5910):

so it's none of them for #3

jimthompson5910 (jim_thompson5910):

how many chose biking?

OpenStudy (anonymous):

8 out of 20

jimthompson5910 (jim_thompson5910):

no just focus on the biking

OpenStudy (anonymous):

3 girls out of 20 students

jimthompson5910 (jim_thompson5910):

ignore the 20 for now, all we care about are the people who chose bike

jimthompson5910 (jim_thompson5910):

there are 5 guys and 3 girls total: 5+3 = 8 So there are 8 people total who chose bike

jimthompson5910 (jim_thompson5910):

so this means (# of girls who chose bike)/(# of people total who picked bike) = 3/8 = 0.375 = 37.5%

OpenStudy (anonymous):

okay thank you for explaining, again!

jimthompson5910 (jim_thompson5910):

yw